OpenStudy (anonymous):

I BET NO ONE CAN SOLVE THIS ! Nick opened a savings account 11 years ago with a deposit of $3,772.59. The account has an interest rate of 5.8% compounded twice a year. How much interest has Nick earned?

OpenStudy (anonymous):

savings account with 5.8 interest twice a year? dang i need to switch to his bank

OpenStudy (anonymous):

A=3,772.59(1+0.058/2)^(2*11) <--- what I did A=3,772.59(1.029)^(22) A=3,772.59(1.8756) A=7,075.869804 and its not right .. what am I doing wrong ?